# Measurement-based Fast Quantum State Stabilization with Deep Reinforcement Learning

### v1

@@ Abstract
The stabilization of quantum states is a fundamental problem for realizing various quantum technologies. Measurement-based-feedback strategies have demonstrated powerful performance, and the construction of quantum control signals using measurement information has attracted great interest. However, the interaction between quantum systems and the environment is inevitable, especially when measurements are introduced, which leads to decoherence. To mitigate decoherence, it is desirable to stabilize quantum systems faster, thereby reducing the time of interaction with the environment. In this paper, we utilize information obtained from measurement and apply deep reinforcement learning (DRL) algorithms, without explicitly constructing specific complex measurement-control mappings, to rapidly drive random initial quantum state to the target state. The proposed DRL algorithm has the ability to speed up the convergence to a target state, which shortens the interaction between quantum systems and their environments to protect coherence. Simulations are performed on two-qubit and three-qubit systems, and the results show that our algorithm can successfully stabilize random initial quantum system to the target entangled state, with a convergence time faster than traditional methods such as Lyapunov feedback control. Moreover, it exhibits robustness against imperfect measurements and delays in system evolution.
{IEEEkeywords}
quantum state stabilization, feedback control, learning control, deep reinforcement learning (DRL)
1 Introduction
[…]
@@ 1 Introduction
Recently, quantum learning control, first introduced in [9], has proven potential in addressing various quantum control problems. Its popularity has grown with the incorporation of additional machine learning (ML) algorithms that exhibit excellent optimization performance and promising outcomes. Quantum learning control concerns to apply proper ML algorithms as promising tools for improving quantum system performance [10, 9, 11, 12, 13], and can offer robust solutions for developing effective quantum control and estimation methods [10]. For instance, the utilization of gradient algorithms in quantum learning control has demonstrated its significance in addressing quantum robust control problems [14]. Another example involves a widely used class of algorithms, evolutionary algorithms (EAs), which gains attention in learning control due to their ability to avoid local optima and their independence from gradient information[1]. Nonetheless, the real-time nature and randomness of measurement feedback control pose challenges, expanding the decision space significantly due to randomness [15]. This randomness makes it almost impossible to reproduce the same measurement trajectory, bringing challenges for EAs to apply control policies from certain sample trajectories to entirely different ones.
Deep reinforcement learning (DRL) emerges as a promising solution to tackle these challenges. Since the introduction of the first DRL algorithm, deep Q-network (DQN) [16], in 2013, and the groundbreaking success of AlphaGo [17] defeating the world champion in the game of Go in 2016, DRL has garnered significant attention. As a ML methodology, DRL employs trial-and-error learning to train intelligent agents to make improved decisions in their environment, aiming to maximize cumulative rewards. The primary step involves the DRL agent engaging with the environment, gathering data and enhancing its policy to optimize decision-making across all states. The process continues iteratively, providing “better” control signals to achieve predefined control objectives [1]. DRL thus has found successful applications in quantum control fields to achieve different control objectives, which include stabilizing quantum states [18, 19, 20, 21, 22], optimizing high-fidelity quantum gates [23], and cooling a quantum harmonic oscillator [24]. An intriguing avenue of research using DRL in quantum measurement feedback control involves employing measurement results as training data to develop control strategies for quantum systems. For instance, in the context of cooling the system of a particle in a potential well to their ground state, position measurements are employed, and the measured current is directly used as feedback information to train the DRL agent, resulting in a relatively high-fidelity final state[25]. Recent advancements have also proposed the utilization of the density matrix of the quantum state as input for training DRL agents. These agents use fidelity as a reward function, allowing them to generate simple linear control signals for preparing and stabilizing Fock states of the cavity [20, 21].
Our study is motivated by the application of suitable DRL algorithms within feedback loops to exploit information obtained from measurements, thereby achieving predefined objectives. This approach holds the potential to enhance feedback control schemes, leading to a notable reduction in the convergence time to reach target states and exhibiting robustness in the face of uncertainties. In our study, we aim to devise a feedback control scheme based on DRL algorithms to enhance state stabilization, focusing primarily on entangled quantum states. These states present unique challenges due to their geometric symmetries and multi-qubit nature.
@@
To achieve the objectives, we exploit the information derived from quantum measurement as the input signal to train our DRL agent. The agent actively interacts with the quantum system, making control decisions based on the received input. We design a generalized reward function that quantifies the similarity between the current quantum state and the desired target state. This incentivizes the DRL agent to iteratively learn control strategies that maximize rewards, ultimately leading to more effective control strategies for stabilizing entangled quantum states. Our work contributes to the field of quantum information processing and quantum computation, shedding light on the potential of DRL in addressing complex quantum control challenges.
@@
    A DRL algorithm is proposed to achieve the stabilization of given entangled states. We design an effective and versatile reward function based on the distance between the system state and the target state, allowing flexible parameter adjustment for different objectives to enhance the performance of the DRL agent.
@@
    We compare the proposed DRL-based control strategy with the Lyapunov method for Bell state and GHZ state preparation through numerical simulations. The DRL-based control fields achieve a faster stabilization for both target states, which effectively reduces the noise generated during system-environment interactions.
@@
The following is the organization of this paper. Section 2 briefly introduces the stochastic master equation for quantum systems under continuous weak measurements. Section 3 explains in detail the logic and implementation behind DRL. Section 4 gives some details of the implementation of DRL in the quantum measurement feedback control. Numerical results are given in Section 5. Section 6 is the conclusion.
2 Quantum System Dynamics
For a quantum system, its state can be represented by a density matrix $\rho$ defined in the Hilbert space $\mathbb{H}$. This density matrix exhibits essential properties: Hermitian ($\rho=\rho^{\dagger}$), trace unity (${{\rm Tr}}(\rho)=1$), and positive semi-definite ($\rho\geq 0$). The dynamics of the quantum system may be observed through continuous weak measurements, enabling us to acquire pertinent measurement information for the design of an appropriate feedback controller. The evolution of the quantum trajectory can be described using the stochastic master equation (SME) [15, 26]:
@@ 2 Quantum System Dynamics
${{\color[rgb]{0,0,0}dy(t)=\sqrt{\eta_{c}\kappa_{c}}{\rm Tr}[(c+c^{\dagger})\rho_{t}]dt+dW.}}$ (3)
With the measurement result $y_{t}$, statistical information on the standard Wiener process $dW$ can be collected from Eq. (3). Utilizing Eq. (1), an estimate of the system state can be obtained and utilized to construct a feedback controller.
@@
3 Deep Reinforcement Learning
Abstracting a real-world problem into a Markov decision process (MDP) serves as the foundational step in applying DRL [29]. MDP provides a formal framework for modeling the interaction between an agent and its environment (quantum systems in this work), offering a structured specification of the agent’s decision-making problem. The environment is abstracted with essential elements such as states, actions, rewards, and more, effectively describing the dynamic interaction between the agent and the external world. The agent is a pivotal component of DRL, representing a learning entity or decision-maker that, through interactions with the environment, learns to take actions to achieve objectives and continually refines its decision strategies to enhance the effectiveness of its actions. This process of agent-environment interaction and learning constitutes the core mechanism through which DRL efficiently tackles real-world challenges and achieves desirable outcomes.
@@ 3 Deep Reinforcement Learning
  • $\mathcal{A}$ signifies the set of actions, incorporating the actions $a_{t}\in\mathcal{A}$ that the agent can undertake at each time step. In this context, the actions correspond to the control signals $u_{j}(t)$ defined in Eq. (1), with values ranging from any bounded control strength, for example, $[0,1]$ in this paper.
  • $R$ denotes the reward function, where the immediate reward $r_{t}$ at time $t$ can be influenced by the current state $\rho_{t}$, the action $a_{t}$, and the subsequent state $\rho_{t+1}$ [25]:
    $r_{t}=R(\rho_{t},a_{t},\rho_{t+1}).$ (4)
    This paper considers the task of stabilizing the current state to the target state, thus the reward $r_{t}$ can be simplified as
@@
    At each time step, the agent receives an immediate reward based on the current state $\rho_{t}$. In this study, our reward function is associated with the trace-based distance $D_{\rho_{t}}$ between the current state $\rho_{t}$ and the target state $\rho_{d}$:
    $D_{\rho_{t}}\triangleq 1-{\rm Tr}(\rho_{d}\rho_{t}).$ (6)
    For common entangled states such as Bell states and GHZ states, when the distance $D_{\rho_{t}}=0$, the system state is stabilized at the target state. The agent’s objective is to learn an optimal policy by maximizing the long-term cumulative reward.
  • $P(\rho_{t+1}|\rho_{t},a_{t})$ is the state transition function. It indicates how the environment transitions to the next state $\rho_{t+1}$ after taking action $a_{t}$ in the current state $\rho_{t}$. It is consistent with the stochastic evolution of the quantum system described in Eq. (1), capturing the dynamic nature and randomness of the environment.
  • $\gamma$ is the discount factor, which strikes a balance between the significance of immediate rewards and future rewards. It determines the emphasis placed on future rewards, influencing the agent’s decision-making process in a long-term perspective.
@@
An MDP is a time-dependent and ongoing process, involving continuous interaction between the agent and the environment. The interaction can be illustrated as shown in Figure 1. For simplicity, we consider the interaction between the agent and the environment as a discrete time series (e.g., $t=0,1,2,3,\cdots$). Starting from $t=0$ with the known initial state $\rho_{0}$, the agent selects an action $a_{0}$ based on the density matrix of $\rho_{0}$ and applies it to the environment. Subsequently, the environment undergoes a state transition according to the state transition function $P(\rho_{t+1}|\rho_{t},a_{t})$, resulting in the next state $\rho_{1}$, and provides an immediate reward $r_{1}$ based on the reward function $R$. The environment also utilizes a classical computer to solve the SME (1) to estimate the density matrix of $\rho_{1}$, which is then fed back to the agent. This process is iterated until completion. Therefore, the MDP and the agent jointly generate a sequence or trajectory as follows:
$\rho_{0},r_{0},a_{0},\rho_{1},r_{1},a_{1},\rho_{2},r_{2},a_{2},\rho_{3},r_{3},\cdots$
The function that selects an action $a$ from the set of actions $\mathcal{A}$ based on the current state $\rho$ is referred to as a policy $\pi(a|\rho)$. It simulates a mapping function between state inputs and action outputs. The objective of the MDP is to find the policy $\pi(a|\rho)$ that allows the agent to make optimal decisions, effectively maximizing long-term cumulative rewards. It is important to note that the resulting state $\rho$ may not be a valid quantum state due to the inherent randomness in measurements and cumulative errors in solving the SME equation using classical computers [30]. For example, the state matrix may contain non-physical (negative) eigenvalues. To address this issue, a suitable approach is to check the eigenvalues of the estimated matrix at each step. When negative values are encountered, the state should be projected back onto a valid physical state. This can be achieved by finding the closest density matrix under the 2-norm [31]. This approximation ensures that a non-physical density matrix is transformed into the most probable positive semi-definite quantum state with a trace equal to $1$.
DRL has proven to be a highly effective approach for solving MDP problems by using deep neural networks to approximate the policy function $\pi(a|\rho)$ of the agent. This enables DRL to solve high-dimensional and complex decision-making problems efficiently. DRL can automatically extract meaningful features from raw inputs and undergo end-to-end training within the reinforcement learning framework [32].
Existing DRL methods can be categorized into three main types: value-based methods, policy-based methods, and actor-critic methods. The value-based approach aims to learn an action-value function $Q(\rho,a)$ which represents the value of taking an action $a$ in state $\rho$. The agent typically uses an $\epsilon$-greedy approach to select the appropriate action in a given state, resulting in a policy function $\pi(a|\rho)$ that guides the agent’s decision; there is no explicit policy in the learning process. One well-known algorithm in this category is DQN [16], which uses a deep neural network to approximate the value function and iteratively updates the network to approximate the optimal value function, allowing for optimal decision-making.
Policy-based methods directly learn the policy function $\pi(a|\rho)$, which represents the probability distribution of selecting action $a$ given state $\rho$. This approach is advantageous for handling continuous action spaces. A common algorithm in this category is policy gradient (PG) [33], which maximizes cumulative rewards using gradient ascent to improve policy performance.
Actor-critic methods are a hybrid approach that simultaneously learns the value function and policy function. The policy function, known as the “actor”, interacts with the environment by selecting actions based on the current state, while the value function, known as the “critic”, evaluates the value of actions. This method leverages the experience of the actor to enhance the critic’s evaluation and, in turn, improves the actor’s policy through the critic’s evaluation [34].
Overall, these three categories of DRL methods offer distinct approaches to address various challenges and provide a wide range of techniques for learning optimal decision-making policies. In the following section, we delve into the fundamental principles of the DRL, leading to the application of the highly effective actor-critic style proximal policy optimization (PPO) algorithm [35] in this paper.
4 Applying DRL to Quantum Measurement-based Feedback Control
@@
4.1 Environment: States and Actions
The two most basic elements of the environment are states and actions. The prior step in applying DRL is to convert the quantum states into a form that can be recognized by DRL. To facilitate agent training and testing, the quantum density matrix is sequentially flattened into a list that contains both the real and imaginary parts (vectorization). For example, the density matrix $\rho=\begin{bmatrix}\alpha_{1}+\beta_{1}i&\alpha_{2}+\beta_{2}i\\ \alpha_{3}+\beta_{3}i&\alpha_{4}+\beta_{4}i\end{bmatrix}$ of a single qubit is transformed into $\rho:=\begin{bmatrix}\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4}\end{bmatrix}^{T}$. Additionally, in order to better diagnose the learned policy, the current evolution time is also included as one of the observed states during agent training. By having access to the current time, the agent is more likely to exhibit improved performance[36]. Therefore, the observation received by the agent is in the form of $\rho(t):=\begin{bmatrix}\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4},t\end{bmatrix}^{T}$.
@@ 4.1 Environment: States and Actions
As for actions, the policy specifies the control signals corresponding to control Hamiltonians. For example, if the system in Eq. (1) has two control Hamiltonians $H_{1}$ and $H_{2}$, the actions for each time step are in the form of $a:=\begin{bmatrix}u_{1},u_{2}\end{bmatrix}^{T}$.
Guided by the policy $\pi_{\theta}(a|\rho)$, the agent stochastically selects an action $a_{t}$ for an environmental state $\rho_{t}$. The environment then transitions to a new state $\rho_{t+1}$ and provides an immediate reward $r_{t}$. Subsequently, the agent selects a new action $a_{t+1}$ based on the updated state $\rho_{t+1}$, and this cycle continues until a predefined termination criterion is met, such as reaching a maximum step count, achieving a goal, or exceeding an acceptable error threshold. This iterative process is termed an episode (or sequence) $\tau=\{\rho_{0},a_{0},\rho_{1},a_{1},\dots,\rho_{m}\}$, where $m$ denotes the final step of the episode, and $\tau\in\mathbb{D}=\{\tau_{n}\}_{n=1,2,\dots,N}$, with $N$ denoting the total number of possible sequences. Without any confusion, we use $\tau$ in the following to represent one episode. The agent’s primary objective is to maximize a measure of expected cumulative reward across all potential episodes.
4.2 Reward
The immediate reward $r_{t}$ at each step is closely related to the distance $D_{\rho_{t}}$ in Eq. (6). We expect a positive reward as the distance approaches 00. To encourage the system to reach the target state, inspired by the inverse proportional functional form, we define a reward function (IPF Reward) with different behaviors based on the distance to the goal state as

### v2

# Fast State Stabilization using Deep Reinforcement Learning for Measurement-based Quantum Feedback Control

@@ Abstract
The stabilization of quantum states is a fundamental problem for realizing various quantum technologies. Measurement-based-feedback strategies have demonstrated powerful performance, and the construction of quantum control signals using measurement information has attracted great interest. However, the interaction between quantum systems and the environment is inevitable, especially when measurements are introduced, which leads to decoherence. To mitigate decoherence, it is desirable to stabilize quantum systems faster, thereby reducing the time of interaction with the environment. In this paper, we utilize information obtained from measurement and apply deep reinforcement learning (DRL) algorithms, without explicitly constructing specific complex measurement-control mappings, to rapidly drive random initial quantum state to the target state. The proposed DRL algorithm has the ability to speed up the convergence to a target state, which shortens the interaction between quantum systems and their environments to protect coherence. Simulations are performed on two-qubit and three-qubit systems, and the results show that our algorithm can successfully stabilize random initial quantum system to the target entangled state, with a convergence time faster than traditional methods such as Lyapunov feedback control and several DRL algorithms with different reward functions. Moreover, it exhibits robustness against imperfect measurements and delays in system evolution.
Index Terms:
deep reinforcement learning (DRL), feedback control, learning control, quantum state stabilization
I Introduction
@@
Recently, quantum learning control, first introduced in [9], has proven potential in addressing various quantum control problems. Its popularity has grown with the incorporation of additional machine learning (ML) algorithms that exhibit excellent optimization performance and promising outcomes. Quantum learning control concerns to apply proper ML algorithms as tools for improving quantum system performance [10, 9, 11, 12, 13], and can offer robust solutions for developing effective quantum control and estimation methods [10]. For instance, the utilization of gradient algorithms in quantum learning control has demonstrated its significance in addressing quantum robust control problems [14]. Another example involves a widely used class of algorithms, evolutionary algorithms (EAs), which gains attention in learning control due to their ability to avoid local optima and their independence from gradient information[1]. Nonetheless, the real-time nature and randomness of measurement feedback control pose challenges, expanding the decision space significantly due to randomness [15]. This randomness makes it almost impossible to reproduce the same measurement trajectory, bringing challenges for EAs to apply control policies from certain sample trajectories to entirely different ones.
Our study is motivated by the application of suitable deep reinforcement learning (DRL) algorithms within feedback loops to exploit information obtained from measurements, thereby achieving predefined objectives. This approach holds the potential to enhance feedback control schemes, leading to a reduction in the convergence time to reach target states and exhibiting robustness in the face of uncertainties. RL techniques have been applied for target state preparation in many-body quantum systems of interacting qubits, where a sequence of unitaries was applied and a measurement was implemented at the final time to provide reward information to the agent [16]. The similar idea was then utilized in harmonic oscillator systems, to achieve different target state preparations, through an ancilla qubit in [17], where a final reward measurement (POVM) was carefully designed to provide reward information to the agent for different tasks of state preparation. In recent years, DRL (RL) approaches also started to play an important role in quantum algorithms, such as QAOA, for ground state preparation in different quantum systems [18, 19]. The similar state preparation problem has also been considered in [20], where the system is a double-well system and the reward is a unique function of the measurement results. In [21, 22], DRL-based approaches were employed for the preparation of cavity Fock state superpositions using fidelity-based reward functions, with system states as the training information under continuous weak measurement. In this study, we aim to devise a feedback control scheme based on DRL algorithms to enhance state stabilization, such as Bell states and GHZ states, for multi-qubit systems under continuous weak measurement. The designed algorithm can be applied to multi-qubit systems and provides a high fidelity and faster convergence to a given target state.
@@
To achieve the objectives, we exploit the information derived from quantum measurement as the input signal to train our DRL agent. The agent actively interacts with the quantum system, making control decisions based on the received input. We design a generalized reward function that quantifies the similarity between the current quantum state and the desired target state. This incentivizes the DRL agent to iteratively learn control strategies that maximize rewards, ultimately leading to more effective control strategies for stabilizing entangled quantum states. Our work shows the potential of DRL in solving complex quantum control challenges, contributing to the fields of quantum information processing and quantum computation.
@@
    A DRL algorithm is proposed to achieve the stabilization of given entangled states in multi-qubit systems under continuous measurement. We design an effective and versatile reward function based on the distance between the system state and the target state, allowing flexible parameter adjustment for different objectives to enhance the performance of the DRL agent.
@@
    We compare the proposed DRL-based control strategy with the Lyapunov method and other DRL methods, for state preparation through numerical simulations. Our DRL method achieves a faster stabilization for both target states, which effectively reduces the noise generated during system-environment interactions.
@@
  4. 4.
    We conduct ablation studies to showcase the superiority of the proposed reward function. By comparing several commonly used reward function designs, our reward function shows better performance in stabilizing the target states.
The following is the organization of this paper. Section II briefly introduces the stochastic master equation for quantum systems under continuous weak measurements. Section III explains in detail the logic and implementation behind DRL. Section IV gives some details of the implementation of DRL in the quantum measurement feedback control. Numerical results are given in Section V. In Section VI, the performance of the proposed algorithm is analyzed through ablation studies and comparisons with other related methods. Section VII is the conclusion.
II Quantum System Dynamics
For a quantum system, its state can be represented by a density matrix $\rho$ defined in the Hilbert space $\mathbb{H}$. This density matrix exhibits essential properties: Hermitian ($\rho=\rho^{\dagger}$), trace unity (${\rm{Tr}}(\rho)=1$), and positive semi-definite ($\rho\geq 0$). The dynamics of the quantum system may be observed through continuous weak measurements, enabling us to acquire pertinent measurement information for the design of an appropriate feedback controller. The evolution of the quantum trajectory can be described by the stochastic master equation (SME) [15, 23]:
@@ II Quantum System Dynamics
${dy(t)=\sqrt{\eta_{c}\kappa_{c}}\rm{Tr}[(c+c^{\dagger})\rho_{t}]dt+dW.}$ (3)
With the measurement result $y_{t}$, information on the standard Wiener process $dW$ can be collected from (3). Utilizing (1), an estimate of the system state can be obtained and utilized to construct a feedback controller.
@@
III Deep Reinforcement Learning
Abstracting a real-world problem into a Markov decision process (MDP) serves as the foundational step in applying DRL [26]. MDP provides a formal framework for modeling the interaction between an agent and its environment (quantum systems in this work), offering a structured specification of the agent’s decision-making problem. The environment is abstracted with essential elements such as states, actions, rewards, and more. The agent is a pivotal component of DRL, representing a learning entity or decision-maker that, through interactions with the environment, learns to take actions to achieve objectives and continually refines its decision strategies to enhance the effectiveness of its actions. This process of agent-environment interaction and learning constitutes the core mechanism through which DRL efficiently tackles real-world challenges and achieves desirable outcomes.
@@ III Deep Reinforcement Learning
  • $\mathcal{A}$ signifies the set of actions, incorporating the actions $a_{t}\in\mathcal{A}$ that the agent can undertake at each time step. In this context, the actions correspond to the control signals $u_{j}(t)$ defined in (1), with values ranging from any bounded control strength, for example, $u_{j}(t)\in[-1,1]$ in this paper.
  • $R$ denotes the reward function. This paper considers the task of stabilizing the current state to the target state, thus the immediate reward $r_{t}$ can be simplified as
@@
    In this study, the reward function is defined using the trace-based distance $D_{\rho_{t}}$:
    $D_{\rho_{t}}\rm Triangleq1-\rm Tr(\rho_{d}\rho_{t}),$ (5)
    which quantifies the difference between the current state $\rho_{t}$ and the target state $\rho_{d}$. When $D_{\rho_{t}}=0$, the system state has stabilized at the target state $\rho_{d}$.
  • $P(\rho_{t+1}|\rho_{t},a_{t})$ is the state transition function. It indicates how the environment transitions to the next state $\rho_{t+1}$ after taking action $a_{t}$ in the current state $\rho_{t}$. It is consistent with the stochastic evolution of the quantum system described in (1).
  • $\gamma$ is the discount factor, which determines the emphasis placed on future rewards, influencing the agent’s decision-making process in a long-term perspective.
@@
An MDP is a time-dependent and ongoing process, involving continuous interaction between the agent and the environment. The interaction can be illustrated as shown in Fig. 1. For simplicity, we consider the interaction between the agent and the environment as a discrete time series (e.g., $t=0,1,2,3,\cdots$). Starting from $t=0$ with the known initial state $\rho_{0}$, the agent selects an action $a_{0}$ based on $\rho_{0}$ and applies it to the environment. Subsequently, the environment undergoes a state transition according to the state transition function $P(\rho_{t+1}|\rho_{t},a_{t})$, resulting in the next state $\rho_{1}$, and provides an immediate reward $r_{1}$ based on the reward function $R$. The environment also utilizes a classical computer to solve the SME (1) to estimate the density matrix of $\rho_{1}$, which is then fed back to the agent. This process is iterated until completion. Therefore, the MDP and the agent jointly generate a sequence or trajectory as follows:
$\{\rho_{0},r_{0},a_{0}\},\{\rho_{1},r_{1},a_{1}\},\{\rho_{2},r_{2},a_{2}\},\cdots$ (6)
The function that selects an action $a$ from the set of actions $\mathcal{A}$ based on the current state $\rho$ is referred to as a policy $\pi(a|\rho)$. The objective of the MDP is to find the policy $\pi(a|\rho)$ that allows the agent to make optimal decisions, effectively maximizing long-term cumulative rewards. Different methods, including value-based, policy-based and actor-critic-based, have been developed to update the policy in the DRL [27, 28, 29]. In this paper, a highly effective actor-critic style proximal policy optimization (PPO) algorithm [30] will be applied.
IV Applying DRL to Quantum Measurement-based Feedback Control
@@
IV-A Environment: States and Actions
The environment’s state is represented by the quantum system’s density matrix, $\rho$, which contains all the information about the quantum state. To use $\rho$ in DRL, it needs to be converted into a format suitable for neural networks. We achieve this by flattening the density matrix into a vector that includes both its real and imaginary parts. For example, for a single-qubit system with $\rho=\begin{bmatrix}\alpha_{1}+\beta_{1}i&\alpha_{2}+\beta_{2}i\\ \alpha_{3}+\beta_{3}i&\alpha_{4}+\beta_{4}i\end{bmatrix},$ it is converted to: $\rho:=\begin{bmatrix}\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4}\end{bmatrix}^{T}$.This process keeps all the quantum state information and makes it usable for the neural network.
@@ IV-A Environment: States and Actions
The policy function $\pi_{\theta}(a|\rho)$ maps the quantum state’s vectorized form to a set of control actions $a$. These actions correspond to the control signals applied to the control Hamiltonians $H_{1},H_{2},\dots,H_{M}$. For example, if there are two control Hamiltonians, the actions at each time step are given by $a:=\begin{bmatrix}u_{1},u_{2}\end{bmatrix}^{T}$, where $u_{1}$ and $u_{2}$ are the control amplitudes.
The DRL algorithm trains the policy to select actions that stabilize the quantum state. At each time step $t$, the agent uses the policy to choose an action $a_{t}$ based on the state $\rho_{t}$. The environment then updates the quantum state to $\rho_{t+1}$ according to the system dynamics and provides a reward $r_{t}$. This reward encourages the agent to take actions that reduces the difference between $\rho_{t+1}$ and the target state $\rho_{\text{target}}$.
It is important to note that the resulting state $\rho$ may not be a valid quantum state due to the inherent randomness in measurements and cumulative errors in solving the SME using classical computers [31]. For example, the state matrix may contain non-physical (negative) eigenvalues. To address this issue, a suitable approach is to check the eigenvalues of the estimated matrix at each step. When negative values are encountered, the state should be projected back onto a valid physical state. This can be achieved by finding the closest density matrix under the 2-norm [32]. This approximation ensures that a non-physical density matrix is transformed into the most probable positive semi-definite quantum state with a trace equal to $1$.
As mentioned in Section III, the interaction between the agent and the environment forms an episode $\tau=\{\rho_{0},r_{0},a_{0},\rho_{1},r_{1},a_{1},\dots,\rho_{m}\},$ where $m$ is the final step of the episode, $\tau\in\mathbb{D}=\{\tau_{n}\}_{n=1,2,\dots,N}$, with $N$ denoting the total number of possible sequences. Without confusion, we use $\tau$ to represent a single episode. The primary goal of the agent is to maximize the expected cumulative reward across all possible episodes. The construction of the reward function will be discussed in detail in the following section.
IV-B Reward
The design of the reward function is critical in DRL. For instance, [16, 17] proposed sparse reward functions, i.e., providing rewards only at the end of each trajectory by evaluating the final state. While this approach is straightforward and easy to implement, it often suffers from slow learning or even training failure in complex systems due to the sparsity of reward signals. To address this, [21] and [22] introduced a scheme that collects fidelity information at each step and applies a higher-weighted reward for high-fidelity states. However, these methods still face challenges in balancing exploration and exploitation, particularly in systems with high-dimensional state spaces.
In this work, we propose a Partitioned Nonlinear Reward (PNR) function based on the distance $D_{\rho_{t}}$ (5). A lower $D_{\rho_{t}}$ indicates better alignment with the target state. The reward at each time step $r_{t}$ is defined as: